Solve the equation.
x(2x-13)=15
use commas to separate answers as needed.

Answer by mananth(16949)

x(2x-13)=15
..
2x^2-13x=15
...
2x^2-13x-15=0
2x^2-15x+2x-15=0
x(2x-15)+1(2x-15)=0
(2x-15)(x+1)=0
x=15/2,-1
...
m.ananth@hotmail.ca